You are here: Foswiki>Commissioning Web>Logs>26Aug013_Log (28 Aug 2013, JuanGuerra)Edit Attach

UT 20130826 AO#2 DX.TS4

Observer:JCGuerra(LBTO),LDetman(LBTO),DMiller(LBTO),ARiccardi(Arcetri),RBruglio(Arcetri)
Telescope Operator: David(LBTO)
SW Support:
AO Support:
Telescope Support: )
Instrument Support:

Summary

  • Summer shutdown.
  • Calibration of the TS4 at DX side using the 4D.
  • Checking the flattening obtained with the 500x500 measurements resolution.
  • Taking measurements of noise, to measure the convection.
  • Trigger fibers was replaced by a spare.
  • Build the Actuator mapping.
  • Measurements of the FF-matrix for fine tunning.
  • Removed the actuator 501, multiples jumps.

Detail

Star at 8:00 pm

Trigger.

The trigger fibre is changed by the spare and the test of the trigger shows that the system works ok. The original fibre was gave to Joe Kraus and the fibre checks shows that it is ok.

Hexalign calibration.

@9:00pm The hexapod started a new calibration. AdOpt >print,hexapod_opt_calib(ave=5)

hexapod position: x y z Rx Ry Rz -0.19 0.93 4.18 -166 -144 0

Noise measurements

Noise measurements taken during the night. Taking a few burst we can see some convection with a different pattern than yesterday, this could be cause by the High bay hatch that is completely open.Measurements of the noise is taken; 1000 measurements

file: 20130827_041100_noise

The telescope crew started to close the hatch.

file: 20130827_042537_noise, 20130827_045655_noise,20130827_050556_noise

Produce file: 20130827_042537_noise

20130827_060000_noise, 20130827_061521_noise,20130827_065600_noise,20130827_100300_noise

Flattening half resolution

Flattening half resolution

Check of the 500x500 int mat acquire and redux yesterday.

Track.num n.modes rmsWFE(nm) intmat notes
20130827_035840 400 96 20130826_150704 stable
20130827_040002 500 98 20130826_150704 stable
20130827_040226 600 80 20130826_150704 stable
20130827_050422 600 79 20130826_150704 stable

Actuator Mapping

IDL> a='/towerdata/meas/adsec_calib/if_functions/newmodes_20130826/MMmatrix.sav' IDL> make_offload_matrix_proj,a, /ver

then, the index of pixel corresponding to each capsens may be obtained with: q = draw_act_map(reference_mask, act)

FF-MATRIX Tunning

Command used; print,collect_ff_data(/no)

The script to run the collection of the data didn't work. When the routine rested the shell and the load program, the start up process didn't complete correctly. It is seems that some files; corrupted the startup of the mirror.

@6:00 am the problem is solved and the acquisition of the FF-Matrix is started

act#279 is showing a spike like residual in the data. to be excluded

IF modes collection for the KL basis

@ 9:40 the acquisition of the data

i4d_opt_intmat(600,660,avera=11,FORCE_sc=0.5/0.3)

print, i4d_opt_intmat(i*25,i*25+24, ave=25)

Tracknum modes average amp Notes
20130827_154317 300-324 25
20130827_154421 325-349 25
20130827_155106 350-374 25 using the extraforce here produces a RIP!!
20130827_155201 375,399 25
20130827_155408 400-424 25
20130827_155504 425-449 25
20130827_155558 450-474 25
20130827_155654 475-499 25

Stop Observing 9:30am

Meeting to discuss the option to follow for the problem displayed of the internal ring

-- JuanGuerra - 28 Aug 2013
Topic revision: r1 - 28 Aug 2013, JuanGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback